With a bag of 5 blue marbles, 7 green marbles, and 3 red marbles. What is the probability of picking out two blue marbles? Pick them one at a time and do not replace them.

So there are 15 marbles in total (5 + 7 + 3) The probability of the first marble will be the number of blue marbles / total number of marblesTherefore, you do 5/15 = 1/3 Then for the second marble you do the same. But there are only 4 blue marbles left and 14 in total, so you do 4/14.Because these events are AND events. Probability of picking out a blue first time AND second time. You multiply the two probabilities 1/3 * 4/14 = 2/21

CB

Related Maths GCSE answers

All answers ▸

If the hypotenuse of a triangle is 7cm and another side is 4cm, what's the length of the other side? How can I work this out?


Find the zeroes of the quadratic polynomial x^2 + 8x + 15.


Write 36 as a product of its prime factors giving the answer in index form.


Solve the simultaneous equations 3x + 2y = 4, 4x + 5y = 17